# Music Proofreading with RefinPaint: Where and How to Modify Compositions given Context

### v1

@@ 1 Introduction
The human-in-the-loop approach we propose allows for selecting the number of tokens to modify and revise the analysis heatmap at each iteration, as described in the following section. Through experimentation, we confirm RefinPaint’s effectiveness in inpainting and proofreading tasks, demonstrating its utility for enhancing music created by both machines and humans. Finally, we provide a companion page featuring examples [^1] At: https://refinpaint.github.io/ and the code along with the trained models of RefinPaint to ensure reproducibility [^2] At: https://github.com/ta603/RefinPaint.
@@ 2.1 Inpainting model ($\mathcal{I}$)
The inpainting model aims to predict, or fill in, missing parts of a MIDI sequence based on a given mask. We adopt an encoder-decoder architecture for sequence-to-sequence tasks, inspired by the PIA study for music generation [15]. This model involves an encoder converting input data into a latent representation and a decoder predicting the final output.
@@ 2.1.1 Training the Inpainting Model ($\mathcal{I}$)
0:  MIDI dataset $\mathcal{D}$, Inpainting model $\mathcal{I}$
@@
4:     $M_{u}\leftarrow\text{Fragment}(x,t_{1})$
@@
6:     $L\leftarrow\text{ForwardInpaintingModel}(\mathcal{I},x,M_{s})$$\triangleright$ model forward and compute loss
@@ 2.2.1 Training the Feedback model ($\mathcal{F}$)
0:  MIDI dataset $\mathcal{D}$, Inpainting model $\mathcal{I}$, Feedback model $\mathcal{F}$
@@
4:     $M_{u}\leftarrow\text{Fragment}(x,t_{1})$
5:     $M_{s}\leftarrow\text{Random Masking}(M_{s},\gamma(t_{2}))$
6:     $\hat{x}\leftarrow\mathcal{I}(x,M_{u})$
7:     $L\leftarrow\text{ForwardFeedbackModel}(\mathcal{F},\hat{x},M_{u})$$\triangleright$ model forward and compute loss
@@
One ideal way of training $\mathcal{F}$ would involve a vast dataset of computer- or human-generated music compositions and human experts’ revisions for inpainting and proofreading applications. Instead, we propose a more feasible synthetic training strategy. The inpainting model $\mathcal{I}$ generates tokens within the selected fragment of a music piece, $M_{u}$, which we label as ‘Fake’, while we label as ‘Real’ the original unchanged tokens. We utilize these labels to instruct $\mathcal{F}$, following the process illustrated in Figure 4.
@@ 2.3 Generation of MIDI sequences (RefinPaint)
We capitalize on the strengths of the inpainting and feedback models for the iterative MIDI sequence generation. The process begins with a MIDI sequence $x$ introduced by the user, setting the stage for a loop that spans a predetermined number of iterations $T$.
@@
At each iteration $t$, the inpainting model $\mathcal{I}$ generates a new version of the sequence $\hat{x}$, based on the current masked input $x_{m}^{(t)}$. In the human-in-the-loop scenario, the user can then adjust this generated sequence. The feedback model $\mathcal{F}$ evaluates $\hat{x}$ and provides a new mask $M^{(t+1)}$, which the user may also modify. This mask highlights the tokens that are deemed most realistic. The number of selected realistic tokens $k$ follows a decreasing function $\gamma$ of the iteration $t$, which models the increasing confidence in the tokens produced over time. Moreover, we add an extra binary embedding to the encoder input with information about the mask M–the given context–where M changes over iterations.
@@
5:        $M^{(i+1)}\leftarrow\mathcal{F}(\hat{x})$
6:        $x_{m}^{(i+1)}\leftarrow\text{k-realistic tokens}(\hat{x},M^{(i+1)},k)$
@@ 5.3 Listening test of proofreading inpainting
For each experiment, which involved 50%, 30%, and 10% fragments of inpainted content, 15 different annotators evaluated both the first iteration of inpainted content (PIA) and the complete iterative process of RefinPaint (Ours) for ten iterations. Participants were exposed to two scenarios, Experiment 1 and Experiment 2: one from the PIA model and one from our RefinPaint model. The order in which these pairs were presented was randomized to avoid any bias. Additionally, we provided the original music fragment without the inpainted content for reference. Participants listened to both the PIA and RefinPaint versions before making their evaluations. They were asked to assess the inpainted content’s quality by comparing it to their first draft, focusing specifically on coherence and creativity. To make their choice, participants were given four options to prevent bias: ‘Experiment 1,’ ‘Maybe Experiment 1,’ ‘Maybe Experiment 2,’ and ‘Experiment 2’.

### v2

@@ 1 Introduction
The human-in-the-loop approach we propose allows for selecting the number of tokens to modify and revise the analysis heatmap at each iteration, as described in the following section. Through experimentation, we confirm RefinPaint’s effectiveness in inpainting and proofreading tasks, demonstrating its utility for enhancing music created by both machines and humans. Finally, we provide a companion page featuring examples [^1] At: https://refinpaint.github.io/ and the code along with the trained models of RefinPaint for reproducibility [^2] At: https://github.com/ta603/RefinPaint.
@@ 2.1 Inpainting model ($\mathcal{I}$)
The inpainting model aims to predict, or fill in, missing parts of a MIDI sequence based on a given mask. We adopt an encoder-decoder architecture for sequence-to-sequence tasks, as shown in Figure 4, inspired by the PIA study for music generation [15]. This model involves an encoder converting input data into a latent representation and a decoder predicting the final output.
@@ 2.1.1 Training the Inpainting Model ($\mathcal{I}$)
0:  MIDI dataset $\mathcal{D}$, Inpainting model $\mathcal{I}$, ground-truth label $y$
@@
4:     $M_{u}\leftarrow\text{Fragment}(t_{1})$
@@
6:     $L\leftarrow\text{CE}(\mathcal{I}(x\odot M_{s}),y\odot M_{s})$
@@ 2.2.1 Training the Feedback model ($\mathcal{F}$)
0:  MIDI dataset $\mathcal{D}$, Inpainting model $\mathcal{I}$, Feedback model $\mathcal{F}$, Real or fake label creation function $\text{Label}(\cdot)$
@@
4:     $M_{u}\leftarrow\text{Fragment}(t_{1})$
5:     $M_{s}\leftarrow\text{RandomMasking}(M_{u},\gamma(t_{2}))$
6:     $\hat{x}\leftarrow\mathcal{I}(x\odot M_{s})$
7:     $L\leftarrow\text{BCE}(\mathcal{F}(\hat{x},M_{u})\odot M_{u},\text{Label}(M_{s})\odot M_{u})$
@@
One ideal way of training $\mathcal{F}$ would involve a vast dataset of computer- or human-generated music compositions and human experts’ revisions for inpainting and proofreading applications. Instead, we propose a more feasible synthetic training strategy, described in Algorithm 2. The inpainting model $\mathcal{I}$ generates tokens within the selected fragment of a music piece, $M_{u}$, which we label as ‘Fake’, while we label as ‘Real’ the original unchanged tokens. We utilize these labels to instruct $\mathcal{F}$, following the process illustrated in Figure 4.
@@ 2.3 Generation of MIDI sequences (RefinPaint)
We capitalize on the strengths of the inpainting and feedback models for the iterative MIDI sequence generation. The process shown in Figure 4 begins with a MIDI sequence $x$ introduced by the user, setting the stage for a loop that spans a predetermined number of iterations $T$.
@@
In the proposed Algorithm 3, at each iteration $t$, the inpainting model $\mathcal{I}$ generates a new version of the sequence $\hat{x}$, based on the current masked input $x_{m}^{(t)}$. In the human-in-the-loop scenario, the user can then adjust this generated sequence. The feedback model $\mathcal{F}$ evaluates $\hat{x}$ and provides a new mask $M^{(t+1)}$, which the user may also modify. This mask highlights the tokens that are deemed most realistic. The number of selected realistic tokens $k$ follows a decreasing function $\gamma$ of the iteration $t$, which models the increasing confidence in the tokens produced over time. Moreover, we add an extra binary embedding to the encoder input with information about the mask M–the given context–where M changes over iterations.
@@
5:        $M^{(i+1)}\leftarrow\text{CreateMask}(\mathcal{F}(\hat{x},M_{u}))$
6:        $x_{m}^{(i+1)}\leftarrow\hat{x}\odot M^{(i+1)}$ with $k$-realistic tokens
@@ 5.3 Listening test of proofreading inpainting
For each experiment, which involved 50%, 30%, and 10% fragments of inpainted content, 15 different annotators evaluated both the first iteration of inpainted content (PIA) and the complete iterative process of RefinPaint (Ours) for ten iterations. Participants were exposed to two scenarios, Experiment 1 and Experiment 2: one from the PIA model and one from our RefinPaint model. The order in which these pairs were presented was randomized to avoid any bias. Additionally, we provided the original music fragment without the inpainted content for reference. Participants listened to both the PIA and RefinPaint versions before making their evaluations. They were asked to assess the inpainted content’s quality by comparing it to the original fragment, focusing specifically on coherence and creativity. To make their choice, participants were given four options to prevent bias: ‘Experiment 1,’ ‘Maybe Experiment 1,’ ‘Maybe Experiment 2,’ and ‘Experiment 2’.